# Safe and Generalized end-to-end Autonomous Driving System with Reinforcement Learning and Demonstrations

### v3

@@ Abstract
An intelligent driving system should be capable of dynamically formulating appropriate driving strategies based on the current environment and vehicle status, while ensuring the security and reliability of the system. However, existing methods based on reinforcement learning and imitation learning suffer from low safety, poor generalization, and inefficient sampling. Additionally, they cannot accurately predict future driving trajectories, and the accurate prediction of future driving trajectories is a precondition for making optimal decisions. To solve these problems, in this paper, we introduce a Safe and Generalized end-to-end Autonomous Driving System (SGADS) for complex and various scenarios. Our SGADS incorporates variational inference with normalizing flows, enabling the intelligent vehicle to accurately predict future driving trajectories. Moreover, we propose the formulation of robust safety constraints. Furthermore, we combine reinforcement learning with demonstrations to augment search process of the agent. The experimental results demonstrate that our SGADS can significantly improve safety performance, exhibit strong generalization, and enhance the training efficiency of intelligent vehicles in complex urban scenarios compared to existing methods.
Figure 1: Overview of the safe and generalized end-to-end autonomous driving with reinforcement learning and demonstrations
@@ 1 Introduction
An intelligent autonomous driving systems must be able to handle complex road geometry and topology, complex multi-agent interactions with dense surrounding dynamic objects, and accurately follow the planning and obstacle avoidance. Current, autonomous driving systems in industry are mainly using a highly modularized hand-engineered approach, for example, perception, localization, behavior prediction, decision making and motion control, etc. (Thrun et al., 2006) and (Urmson et al., 2008). Particularly, the autonomous driving decision making systems are focusing on the non-learning model-based methods, which often requires to manually design a driving policy(González et al., 2015) and (Paden et al., 2016). However, the manually designed policy could have two several weaknesses: 1) Accuracy. The driving policy of human heuristics and pre-training model can be suboptimal, which will lead to either conservative or aggressive driving policies. 2) Generality. For different scenarios and complicated tasks, we might need to be redesigned the model policy manually for each new scenario.
@@
To solve those problems, existing works such as Imitation Learning (IL) is most popular approach, which can learn a driving policy by collecting the expert driving data(Bojarski et al., 2016), (Codevilla et al., 2018), (Bansal et al., 2018) and (Chen et al., 2019a). However, those methods can suffer from the following shortcomings for imitation learning: (1) High training cost and sample complexity. (2) Conservation of driving policy. Due to the collect driving data from the human expert, which can only learn driving skills that are demonstrated in the datasets. (3) Limitation of driving performance. The performance of imitation learning policy is impossible exceed human-level performance.
@@
What’s more, the driving policy based on Reinforcement Learning (RL) is also popular method in recent years, which can automatically learn and explore without any human expert data in various kinds of different driving cases, and it is possible to have a better performance than imitation learning. (Lillicrap et al., 2015) introduced a continuous control deep RL algorithm that trains a deep neural network policy for autonomous driving on a simulated racing track. (Wolf et al., 2017) used Deep Q-Network to learn to steer an autonomous vehicle to keep in the track in simulation. (Chen et al., 2018) developed a hierarchical deep RL framework to handle driving scenarios with intricate decision-making processes, such as navigating traffic lights. (Kendall et al., 2019) marked the first application of deep RL in real-world autonomous driving, where they trained a deep lane-keeping policy using only a single front-view camera image as input. (Chen et al., 2021) proposed an interpretable deep RL method for end-to-end autonomous driving. (Nehme & Deo, 2023) proposed safe navigation.(Murdoch et al., 2023) propose a partial end-to-end algorithm that decouples the planning and control tasks. (Zhou et al., 2023) proposes a method to identify and protect unreliable decisions of a DRL driving policy. (Zhang et al., 2023) a framework of constrained multi-agent reinforcement learning with a parallel safety shield for CAVs in challenging driving scenarios. (Liu et al., 2022) propose the Scene-Rep Transformer to enhance RL decision-making capabilities. (Huang et al., 2023) introduces a predictive behavior planning framework that learns to predict and evaluate from human driving data. In addition, combination of RL with Demonstrations are also learned more complex tasks such as dexterous manipulator. The techniques outlined in (Theodorou et al., 2010), (Van Hoof et al., 2015) and (Rajeswaran et al., 2017) have proven to be efficient in merging demonstrations and RL for improving learning speed. (Huang et al., 2024) propose an enhanced human in-the-loop reinforcement learning method, while they rely on human expert performance and can only accomplish simple scenario tasks. However, the existing methods also have some weakness: (1) They can only deal with the simple driving scenario tasks. (2)High costs of learning and exploration. (4) The safety and generalization of intelligent vehicles need further improvement. (4) Such these RL frameworks are mostly based on Partially Observable Markov Decision Process(POMDP)(Åström, 1965),(Hollinger & Sukhatme, 2014),(Agha-Mohammadi et al., 2014),(Indelman et al., 2015),(Karl et al., 2016). However, they cannot accurately predict future driving trajectories, and the accurate prediction of future driving trajectories is a precondition for making optimal decisions.
Figure 2: (a) RL Agent learns potential dynamics from past experience datasets of autonomous vehicle. (b) RL Agent predicts driving behavior in an imaginary space. (c) RL Agent interacts with driving environment.
In conclusion, the existing IL and RL methods can not well solve the safety, sampling efficiency, generalization and prediction for future driving trajectories in autonomous driving. To solve above problems, in this paper, we introduce a Safe and Generalized end-to-end Autonomous Driving System (SGADS) for complex and various scenarios. Our SGADS incorporates variational inference with normalizing flows, enabling the intelligent vehicle to accurately predict driving trajectories. Moreover, we design safety constraint for intelligent vehicle. We further propose a behavior cloning as to augment search process of agent. We also contribute a demonstration dataset by having human experts drive using the G29 steering wheel. The experimental result shows our SGADS greatly augments safety performance, exhibits strong generalization in multiple testing maps, and enhances the training efficiency of intelligent vehicles in intricate urban scenarios compared to the existing methods. Moreover, we designed and explored up to 12 different data input types to evalutate their impact on the safety performance of intelligent vehicles. To sum up, our contributions are:
  • We introduces a novel SGADS designed to address the limitations of existing methods in solving safety, sampling efficiency, generalization problems in autonomous driving system.
  • SGADS incorporates variational inference with normalizing flows, enabling the intelligent vehicle to accurately predict future driving trajectories.
  • The experimental result shows that SGADS is capable of greatly enhancing safety, generalization, and training efficiency for intelligent vehicles.
  • We provide a dataset collected from human expert using the G29 steering wheel. Moreover, we have designed and explored 12 inputs to evalutate their influence on the safety performance of intelligent vehicles.
2 Problem formulation
2.1 System Framework
The proposed framework of our SGADS in Figure 1. Firstly, human expert collects the demonstrations offline using the Logitech G29 steering wheel. Then, the expert demonstrations are used as the initial experience replay buffers for training the entire model. Subsequently, a pre-training process is conducted to establish a model with human expert experience that do not update environmental data during training. The resulting model with human expert experience is then utilize to initialize network parameters for RL policy. Additionally, safety constraint is designed for intelligent vehicle, enhancing their safety performance. In addition, we design and explore 12 different types of images as the model inputs as shown in Figure 3 and Section 4.7.
2.2 Problem Formulation
Visual control(Thrun, 1999),(Silver & Veness, 2010), (Bengtsson et al., 2008) can be defined as a POMDP, the goal of the RL agent is to maximize expected cumulative reward $E[\sum_{t=0}^{\infty}\gamma_{t}r_{t}]$,$\gamma$ is the discount factor. Inspired by (Hafner et al., 2019a), (Hafner et al., 2019b), the RL agent in our SGADS enable the intelligent vehicle to predict future driving trajectories, which mainly consists of the following three parts:
(1) RL Agent learns potential dynamics from past experience datasets of autonomous vehicle to predict future rewards from actions and past observations. As shown in Figure 2(a), the entire learning process is a POMDP decision chain, using $p$ to represent prior probability, $q$ to represent posterior probability, agent learns to encode observations $o$ and actions $a$ into compact latent states $s$, and $\hat{o}_{t}$ is reconstructed with probability $q(\hat{o}_{t}|s_{t})$ while $s_{t}$ is determined via probability $p(s_{t}|s_{t-1},a_{t-1},o_{t})$.
(2) RL Agent predicts state values and actions of driving behaviors in an imaginary space, which maximize future value predictions by propagating gradients back through imagined trajectories. As shown in Figure 2(b), RL Agent is in a close latent state space where it can predict state $\hat{v}_{t}$, reward $\hat{r}_{t}$ and action $\hat{a}_{t}$ based on current input $o_{t-1}$ with probabilities $q(\hat{v}_{t},\hat{r}_{t},\hat{a}_{t}|s_{t})$, $p(s_{t}|s_{t-1},\hat{a}_{t-1})$, $q(\hat{a}_{t-1}|s_{t-1})$.
(3) RL Agent encodes the history of the episode to compute the current model state and predict the next action to execute in the environment. As shown in Figure 2(c), RL Agent predicts next action values $\hat{a}_{t+1}$ by encoding historical trajectory information via calculating probabilities $q(\hat{a}_{t+1}|s_{t+1})$, $p(s_{t+1}|s_{t},a_{t},o_{t+1})$.
Figure 3: Multiple types of input images
@@ 3 Methodology
3.1 Model Learning
We introduces a novel SGADS model consists of components can be constructed the PGM of POMDP as follow:
@@ 3.1 Model Learning
Belief inference model: $\quad q_{\psi}(s_{t}|\tau_{t},o_{t})$
Where $p$ is prior distribution, $q$ is posterior distribution, $\tau_{t}=\{o_{1},a_{1},\cdots,o_{t-1},a_{t-1},o_{t},a_{t}\}$. The above of components can be optimized by maximizing the Evidence Lower BOund (ELBO)(Jordan et al., 1998),(De Cao et al., 2020):
$\mathcal{M}_{\text{model}}(\theta,\psi,\omega)=\sum_{t=1}^{T}\big{(}\mathbb{E}_{q(s_{t}\mid o_{\leq t},a_{<t})}[\log p(o_{t}\mid s_{t})+$ (2)
$\log p(r_{t}\mid s_{t})]-\big{.}{\mathbb{E}_{q(s_{t-1}\mid\tau_{t-1},o_{t-1})}}[D_{\rm KL}\big{(}q(s_{t}\mid\tau_{t},o_{t})\|$
$p(s_{t}\mid s_{t-1},a_{t-1})]\big{)}\big{)}$
In the continuous and dynamic space, existing methods based on Normalizing Flows(NF)(Dinh et al., 2016) , (Huang et al., 2018), (Rezende & Mohamed, 2015) can learn more flexible and generalized beliefs. These methods provide a solid foundation for RL agents based on POMDP to accurately predict future driving trajectories. Inspired by (Chen et al., 2022), we can substitute the probability density with Normalizing Flow in the KL-divergence term of Equation 2.
$q_{K}(s_{t}|\tau_{t},o_{t})=\log q_{0}(s_{t}|\tau_{t},o_{t})-\sum_{k=1}^{K}|det\frac{\partial f_{\theta_{k}}}{\partial s_{t,k-1}}|$ (3)
@@
Where $q_{K}=q_{\psi,\theta}$, $p_{K}=q_{\psi,\omega}$. The final inferred belief is obtained by propagating $q_{\psi}(s_{t}|\tau_{t},o_{t})$ through a set of NF mappings denoted $f_{\theta_{K}}\dots f_{\theta_{1}}$ to get a posterior distribution $q_{\psi,\theta}(s_{t}|\tau_{t},o_{t})$. The final prior is obtained by propagating $p_{\psi}(s_{t}|\tau_{t})$ through a set of NF mappings denoted $f_{\omega_{K}}\dots f_{\omega_{1}}$ to get a prior distribution $p_{\psi,\omega}(s_{t}|\tau_{t})$.
3.2 Policy Learning
If the initialization of the policy is poor, the learning process may be very slow and the algorithm will explore states and actions that are irrelevant to the task. To address this problem, we use behavior cloning to provide a policy $\pi^{*}$ via expert demonstrations and then to train a model $\mathcal{M}_{\text{expert}}$ with some expert ability. The loss function for training the model in the first stage is as follows:

### v8

# Efficient and Generalized end-to-end Autonomous Driving System with Latent Deep Reinforcement Learning and Demonstrations

@@ Abstract
An intelligent driving system should dynamically formulate appropriate driving strategies based on the current environment and vehicle status while ensuring system security and reliability. However, methods based on reinforcement learning and imitation learning often suffer from high sample complexity, poor generalization, and low safety. To address these challenges, this paper introduces an efficient and generalized end-to-end autonomous driving system (EGADS) for complex and varied scenarios. The RL agent in our EGADS combines variational inference with normalizing flows, which are independent of distribution assumptions. This combination allows the agent to capture historical information relevant to driving in latent space effectively, thereby significantly reducing sample complexity. Additionally, we enhance safety by formulating robust safety constraints and improve generalization and performance by integrating RL with expert demonstrations. Experimental results demonstrate that, compared to existing methods, EGADS significantly reduces sample complexity, greatly improves safety performance, and exhibits strong generalization capabilities in complex urban scenarios. Particularly, we contributed an expert dataset collected through human expert steering wheel control, specifically using the G29 steering wheel. Our code is available: https://github.com/Mark-zjtang/EGADS?tab=readme-ov-file.
Figure 1: Overview of the efficient and generalized end-to-end autonomous driving system with latent deep reinforcement learning and demonstrations.
@@ 1 Introduction
An intelligent autonomous driving systems must be able to handle complex road geometry and topology, complex multi-agent interactions with dense surrounding dynamic objects, and accurately follow the planning and obstacle avoidance. Current, autonomous driving systems in industry are mainly using a highly modularized hand-engineered approach, for example, perception, localization, behavior prediction, decision making and motion control, etc. thrun2006stanley and urmson2008autonomous . Particularly, the autonomous driving decision making systems are focusing on the non-learning model-based methods, which often requires to manually design a driving policy gonzalez2015review and paden2016survey . However, the manually designed policy could have two several weaknesses: 1) Accuracy. The driving policy of human heuristics and pre-training model can be suboptimal, which will lead to either conservative or aggressive driving policies. 2) Generality. For different scenarios and complicated tasks, we might need to be redesigned the model policy manually for each new scenario.
@@
To solve those problems, existing works such as imitation learning (IL) is most popular approach, which can learn a driving policy by collecting the expert driving data. However, those methods can suffer from the following shortcomings for imitation learning: (1) High training cost and sample complexity. (2) Conservation. Due to the collect driving data from the human expert, which can only learn driving skills that are demonstrated in the datasets. (3) Limitation of driving performance. What’s more, the driving policy based on reinforcement learning (RL) is also popular method in recent years, which can automatically learn and explore without any human expert data in various kinds of different driving cases, and it is possible to have a better performance than imitation learning. However, the existing methods also have some weakness: (1) Existing methods in latent space are based on specific distribution assumptions, whereas distributions in the real world tend to be more flexible, resulting in a failure to learn more precisely about belief values. (2) High costs of learning and exploration. (3) The safety and generalization of intelligent vehicles need further improvement.
Combining the advantages of RL and IL, the demonstration of enhanced RL is not only expected to accelerate the initial learning process, but also gain the potential of experts beyond performance. In this paper, we introduce an efficient and generalized end-to-end autonomous driving system (EGADS) for complex and varied scenarios. The RL agent in our EGADS combines variational inference with normalizing flows independent of distribution assumptions, allowing it to sufficiently and flexibly capture historical information useful for driving in latent space, thereby significantly reducing sample complexity. In addition, unlike traditional methods that constrain policy actions directly, we integrate safety constraints into the reward function, which allows the agent to consider safety during training, thereby improving its robustness and generalization. To further increase the upper limit of the overall system, we further enhance the RL search process with a dataset of human experts. In particular, we contributed a dataset of human experts to driving by driving the G29 steering wheel. The experimental results show that compared with the existing methods, our EGADS greatly improves the safety performance, shows strong generalization ability in multiple test maps, and significantly reduces the sample complexity. In summary, our contributions are:
  • We present an EGADS framework designed for complex and varied scenarios.
  • The RL agent in EGADS uses variational inference with normalizing flows (NFRL), independent of distribution assumptions, to capture historical driving information in latent space, significantly reducing sample complexity.
  • We incorporate Safety Constraints (SC) directly into the reward function to enable the agent to account for safety considerations during training.
  • By fine-tuning with a small amount of human expert dataset via using the G29 steering wheel, NFRL agents can learn more general driving principles, significantly improving generalization and sample efficiency.
2 Related Work
Imitation learning, which utilizes an efficient supervised learning approach, has gained widespread application in autonomous driving research due to its simplicity and effectiveness. For instance, imitation learning has been employed in end-to-end autonomous driving systems that directly generate control signals from raw sensor inputs LeMero2022ASO ; codevilla2018end ; bansal2018chauffeurnet ; chen2019deep .
@@ 2 Related Work
Deep reinforcement learning (DRL) has demonstrated its strength in addressing complex decision-making and planning problems, leading to a series of breakthroughs in recent years. Researchers have been trying to apply deep RL techniques to the domain of autonomous driving. Lillicarp et.al lillicrap2015continuous introduced a continuous control DRL algorithm that trains a deep neural network policy for autonomous driving on a simulated racing track. Wolf et.al wolf2017learning used Deep Q-Network to learn to steer an autonomous vehicle to keep in the track in simulation. Chen et.al chen2018deep developed a hierarchical DRL framework to handle driving scenarios with intricate decision-making processes, such as navigating traffic lights. Kendall et.al kendall2019learning marked the first application of DRL in real-world autonomous driving, where they trained a deep lane-keeping policy using only a single front-view camera image as input. Chen et.al chen2021interpretable proposed an interpretable DRL method for end-to-end autonomous driving. Nehme et.al nehme2023safe proposed safe navigation. Murdoch et.al murdoch2023partial propose a partial end-to-end algorithm that decouples the planning and control tasks. Zhou et.al zhou2023identify proposes a method to identify and protect unreliable decisions of a DRL driving policy. Zhang et.al zhang2023spatial a framework of constrained multi-agent reinforcement learning with a parallel safety shield for CAVs in challenging driving scenarios. Liu et.al liu2022augmenting propose the Scene-Rep Transformer to enhance RL decision-making capabilities.
By combining the advantages of RL and IL is also a relatively popular method in recent years. The techniques outlined in theodorou2010reinforcement , van2015learning , zhou2023large and zhao2023context have proven to be efficient in merging demonstrations and RL for improving learning speed. Liu et.al Liu2021ImprovedDR propose a novel framework combining RL and expert demonstration to learn a motion control strategy for urban scenarios. Huang et.al huang2023conditional introduces a predictive behavior planning framework that learns to predict and evaluate from human driving data. Huang et.al huang2024human propose an enhanced human in-the-loop reinforcement learning method, while they rely on human expert performance and can only accomplish simple scenario tasks. DPAG rajeswaran2017learning combines RL and imitation learning to solve complex dexterous manipulation problems. Our approach utilizes the potential for reinforcement learning and normalization flows to learn useful information from historical trajectory information, further learning expert demonstrations through DPAG methods.
@@ 3 Methodology
The proposed framework of our EGADS is illustrated in Figure 1. Firstly, human experts collect demonstrations offline using the G29 steering wheel. These expert demonstrations are then utilized as the RL fine-tuning experience replay buffers for training the entire model. Subsequently, a pre-training process is conducted to establish a model with human expert experience that does not update environmental data during training. The resulting model, enriched with human expert experience, is then used to fine-tune the policy for RL agent. Additionally, we have designed safety constraints for the intelligent vehicle, enhancing its safety performance. Furthermore, we explore 12 different types of images as inputs, which can be found in the Appendix  5.
3.1 Preliminaries
We model the control problem as a Partially Observable Markov Decision Process (POMDP), which is defined using the 7-tuple: $(S,A,T,R,\Omega,O,\gamma)$, where $S$ is a set of states, $A$ is a set of actions, $T$ is a set of conditional transition probabilities between states, $R$ is the reward function, $\Omega$ is a set of observations, $O$ is a set of conditional observation probabilities, and $\gamma$ is the discount factor. The goal of the RL agent is to maximize expected cumulative reward $E[\sum_{t=0}^{\infty}\gamma_{t}r_{t}]$. After having taken action $a_{t-1}$ and observing $o_{t}$, an agent needs to update its belief state, which is defined as the probability distribution of the environment state conditioned on all historical information: $b(s_{t})=p(s_{t}\mid\tau_{t},o_{t})$, where $\tau_{t}=\{o_{1},a_{1},\ldots,o_{t-1},a_{t-1}\}$.
Figure 2: (a) RL agent learns potential dynamics from past experience datasets. (b) RL agent predicts driving action in an imaginary space. (c) RL agent interacts with driving environment. Where $o$ is observation, $a$ is action, $s$ is latent state, $\hat{r}_{t}$ is reward, $\hat{o}_{t}$ is reconstructed and $\hat{v}_{t}$ is value.
3.2 Latent dynamic model for autonomous driving
We propose the use of latent variables to solve problems in end-to-end autonomous driving. This potential space is used to encode complex urban driving environments, including visual inputs, spatial features, and road conditions. Historical high-dimensional raw observation data is compressed into this low-dimensional latent space and learned through a sequential latent environment model that learns in conjunction with the maximum entropy RL process. We introduce RL agent model consists of components can be constructed the probabilistic graphical model of POMDP as follow:
@@ 3.2 Latent dynamic model for autonomous driving
where $p$ is prior probability, $q$ is posterior probability, $o$ is observation, $a$ is action, is latent state and $\theta$ is the parameter of the model. Then the world model reconstructs the inputs images from the original sensors, more details can be found in the Appendix  A.7.
3.3 RL agent in the latent space
Visual control thrun1999monte , silver2010monte , bengtsson2008curse can be defined as a POMDP. The traditional components of agents that learn through imagination include dynamics learning, behavior learning, and environment interaction hafner2019dream , hafner2019learning . The RL agent in the latent space in our EGADS mainly includes the following:
(1) RL agent learns potential dynamics from past experience datasets of autonomous vehicle. As shown in Figure 2(a), using $p$ to represent prior probability, $q$ to represent posterior probability, agent learns to encode observation and action into compact latent state, and $\hat{o}_{t}$ is reconstructed with $q(\hat{o}_{t}|s_{t})$ while $s_{t}$ is determined via $p(s_{t}|s_{t-1},a_{t-1},o_{t})$.
(2) RL agent predicts driving action in an imaginary space. As shown in Figure 2(b), RL agent is in a close latent state space where it can predict value $\hat{v}_{t}$, reward $\hat{r}_{t}$ and action $\hat{a}_{t}$ based on current input $o_{t-1}$ with $q(\hat{v}_{t},\hat{r}_{t},\hat{a}_{t}|s_{t})$, $p(s_{t}|s_{t-1},\hat{a}_{t-1})$, $q(\hat{a}_{t-1}|s_{t-1})$.
(3) RL agent interacts with driving environment. As shown in Figure 2(c), RL agent predicts next action values $\hat{a}_{t+1}$ by encoding historical trajectory information via $q(\hat{a}_{t+1}|s_{t+1})$, $p(s_{t+1}|s_{t},a_{t},o_{t+1})$.
3.4 Normalizing Flow for inferred belief
Existing latent RL models in autonomous driving either suffer from the curse of dimensionality or make some assumptions and only learn approximate distributions. This approximation imposes strong limitations and is problematic, whereas distributions in the real world tend to be more flexible. In the continuous and dynamic space, existing methods based on normalizing flows (NF) dinh2016density , huang2018neural , rezende2015variational can learn more flexible and generalized beliefs. These methods provide a solid foundation for RL agents to accurately predict future driving actions. Inspired by chen2022flow , we added a belief inference model: $q_{\theta}(s_{t}|\tau_{t},o_{t})$, where $\theta$ is the parameter of the model. The belief model can be substituted for the probability density with NF in the KL-divergence term of equation 2.
$q_{K}(s_{t}|\tau_{t},o_{t})=\log q_{0}(s_{t}|\tau_{t},o_{t})-\sum_{k=1}^{K}|det\frac{\partial f_{\psi_{k}}}{\partial s_{t,k-1}}|$ (2)
@@ 3.4 Normalizing Flow for inferred belief
where $q_{0}=q_{\theta}$, $q_{K}=q_{\theta,\psi}$, $p_{0}=p_{\theta}$, $p_{K}=p_{\theta,\omega}$, $\psi$ and $\omega$ are the parameters of a series of mapping transformations of the posterior and prior distributions. Where $\tau_{t}=\{o_{1},a_{1},\cdots,o_{t-1},a_{t-1}\}$. The input images $o_{1:t}$ and actions $a_{1:t-1}$ are encoded with $q_{\theta}(s_{t}|\tau_{t},o_{t})$. Then the final inferred belief is obtained by propagating $q_{\theta}(s_{t}|\tau_{t},o_{t})$ through a set of NF mappings denoted $f_{\psi_{K}}\dots f_{\psi_{1}}$ to get a posterior distribution $q_{\theta,\psi}(s_{t}|\tau_{t},o_{t})$. The final prior is obtained by propagating $p_{\theta}(s_{t}|\tau_{t})$ through a set of NF mappings denoted $f_{\omega_{K}}\dots f_{\omega_{1}}$ to get a prior distribution $p_{\theta,\omega}(s_{t}|\tau_{t})$. Where $p_{K}(s_{t}|\tau_{t})=p_{K}(s_{t}|s_{t-1},a_{t-1})$, given the sampled $s_{t-1}$ from $q_{K}(s_{1:t}|\tau_{t},o_{t})$. Finally, our NF inference RL model (NFRL) is optimized by variational inference method, in which the evidence lower bound (ELBO) jordan1998introduction , de2020block is maximized. The loss function is defined as:
$\mathcal{M}_{\text{model}}(\theta,\psi,\omega)=\sum_{t=1}^{T}\big{(}\mathbb{E}_{q(s_{t}\mid o_{\leq t},a_{<t})}[\log p_{\theta}(o_{t}\mid s_{t})+$ (3)
$\log p_{\theta}(r_{t}\mid s_{t})]-\big{.}{\mathbb{E}_{q_{K}(s_{1:T}\mid o_{1:T},a_{1:T-1})}}[D_{\rm KL}\big{(}q_{K}(s_{t}\mid\tau_{t},o_{t})\|$
$p_{K}(s_{t}\mid\tau_{t},o_{t}))]\big{)}$
3.5 Policy optimization
The action model implements the policy and is designed to predict the actions that are likely to be effective in responding to the simulated environment. The value model estimates the expected reward generated by the behavior model at each state $s_{\tau}$.
$a_{\tau}\sim q_{\phi}(a_{\tau}|s_{\tau}),$ $v_{\eta}(s_{\tau})=E_{q_{\phi}}[\sum_{t=t}^{t+H}\gamma^{\tau-t}r_{\tau}]$ (4)
where $\phi$, $\eta$ are the parameters of the approximated policy and value. The obejective of the action model is to use high value estimates to predict action that result in state trajectories
$\mathcal{M}_{\text{actor}}(\phi)=E_{q_{\phi}}(\sum_{\tau=t}^{t+H}V_{\tau}^{\lambda})$ (5)
To update the action and value models, we calculate the value estimate $v_{\eta}(s_{\tau})$ for all states $s_{\tau}$ along the imagined trajectory. $V_{\tau}^{\lambda}$ can be defined as follow:
$V_{\tau}^{\lambda}=(1-\tau)v_{\eta}(s_{\tau+1})+\lambda V_{\tau+1}^{\lambda},\quad\tau\textless{t+H}$ (6)
Then we can train the critic to regress the TD($\lambda$) sutton2018reinforcement target return via a mean squared error loss:
$\mathcal{M}_{\text{critic}}(\eta)=\mathbb{E}\Big{[}\sum_{\tau=t}^{t+H}\frac{1}{2}\left(v_{\eta}\left(s_{\tau}\right)-V_{\tau}^{\lambda}\right)^{2}\Big{]}$ (7)
where $\eta$ denote the parameters of the critic network and $H$ is the prediction horizon. Then the loss function is as follows:
$\underset{\psi,\eta,\phi,\theta,\omega,\eta}{\rm min}\alpha_{0}{\rm\mathcal{M}_{critic}(\eta)}-\alpha_{1}{\rm\mathcal{M}_{actor}(\phi)}-\alpha_{2}{\rm\mathcal{M}_{model}(\theta,\psi,\omega)}$ (8)
we jointly optimize the parameters of model loss $\psi,\theta,\omega$, critic loss $\eta$ and actor loss $\phi$, where $\alpha_{0},\alpha_{1},\alpha_{2}$ are coefficients for different components.
3.6 Safety constraint
In the Gym-Carla benchmark, the reward function proposed by Chen et.al chen2019model is denoted as $f_{1}$. To ensure the intelligent vehicle operates safely and smoothly in complex environments, we incorporated additional safety and robustness constraints into $f_{1}$, denoted as $f_{2}=f_{1}+200r_{ft}+50r_{lt}+2r_{sc}$. $r_{ft}$ is the front time to collision. $r_{lt}$ is lateral time to collision. $r_{sc}$ is the smoothness constraint. For detailed information on $f_{1}$ and $f_{2}$ of reward function, please refer to Appendix  A.4.
(1) Front time to collision. When around vehicles are within the distance of ego vehicle (our agent vehicle) head in our setting, then we can calculate the front time to collision between ego vehicle and around vehicles. Firstly, the speed and steering vector $(s_{\tau},a_{\tau})\in\mathbb{S}$ of the ego vehicle are defined, where $s_{\tau}$ represents the angle vector of vehicle steering and $a_{\tau}$ represents the acceleration vector of the vehicle in local coordinate system. Secondly, two waypoints closest to the current ego vehicle are selected from the given navigation routing as direction vectors $w_{p}$ for the entire route progression, where $\rightarrow$ indicates a vector in world coordinates. The position vectors for both ego vehicle and around vehicles are represented by $(x_{t}^{*},y_{t}^{*})$, respectively. Finally, $\delta_{e}$ and $\delta_{a}$ representing angles between position vectors for ego vehicle and around vehicles with respect to $w_{p}$} are calculated respectively.
${w_{p}}=\big{[}(\frac{{w_{t+1}^{x}}-{w_{t}^{x}}}{2})-({w_{t}^{x}}),(\frac{{w_{t+1}^{y}}-{w_{t}^{y}}}{2})-({w_{t}^{y}})\big{]}$ (9)
$\delta_{e}=\frac{[{v_{t}^{x*}},{v_{t}^{y*}}]\cdot{w_{p}}}{\|{v_{t}^{x*}},{v_{t}^{y*}}\|_{2}\quad\|{w_{p}}\|_{2}},\delta_{a}=\frac{[{v_{t}^{x}},{v_{t}^{y}}]\cdot{w_{p}}}{\|{v_{t}^{x}},{v_{t}^{y}}\|_{2}\quad\|{w_{p}}\|_{2}}$
where, $l$ is the length of the set of waypoints $\mathbb{W}$ stored. The variable $t\in\tau$, and ${w_{t}^{x}}\in\mathbb{W}_{1}$ represents the $x$ coordinate of the first navigation point closest to the intelligent vehicle on its current route at time $t$. Similarly, ${w_{t+1}^{x}}\in\mathbb{W}_{2}$. Furthermore, it is possible to calculate the $F_{ttc}$ as follows:
$F_{ttc}=\frac{\|{x_{t}}-{x_{t}^{*}},{y_{t}}-{y_{t}^{*}}\|_{2}}{\left|\|{v_{t}^{x*}},{v_{t}^{y*}}\|_{2}sin(\delta_{e})-\|{v_{t}^{x}},{v_{t}^{y}}\|_{2}sin(\delta_{a})\right|}$ (10)